The following independent events for New Age Theatre Ltd. during the year ended November 30, 2018, require a transaction journal entry or an adjusting journal entry, or both. The company adjusts its accounts annually.

1. On June 1, 2017, the theatre purchased vehicles for $80,700 cash. The vehicles’ estimated useful life is five years and the company uses straight-line depreciation.
2. The theatre has eight plays each season. This year’s season starts in October 2018 and ends in May 2019 (one play per month). Season tickets sell for $370. On October 1, 460 season tickets were sold for the 2018–2019 season. The theatre credited Unearned Revenue for the full amount received on October 1 and uses a Ticket Revenue account to record revenue earned from season tickets.
3. Supplies on hand amounted to $1,010 at the beginning of the year. On February 16, additional Supplies were purchased for cash at a cost of $2,370. At the end of the year, a physical count showed that supplies on hand amounted to $670.
4. On June 1, 2018, the theatre borrowed $105,300 from the Bank of Montreal at an interest rate of 6%. The principal is to be repaid in one year. The interest is payable on the first day of each following month and was last paid on November 1.
5. The New Age Theatre rents a portion of its facilities for $520 a month to a local dance club that uses the space for rehearsals. On November 2, the dance club’s treasurer made a mistake and accidentally sent a cheque for only $220 for the November rent. (Hint: Use the Unearned Revenue account to record the rent received in advance.) The dance club’s treasurer promised to send a cheque in December for the balance when she returned from vacation. On December 4, the theatre received a $300 cheque for the balance owing from November.
6. The total weekly payroll is $7,300, paid every Monday for employee salaries earned during a seven-day workweek running from Sunday to Saturday. Salaries were last paid (and recorded) on Monday, November 26, and will be paid next on Monday, December 3. November 30 falls on a Friday this year.
7. Upon reviewing its income tax calculations on November 30, the theatre noted that an additional $1,380 of income tax was owed. This additional amount was paid on December 14.

Prepare the journal entry to record the original transaction for items 1, 2, 3, 4, and 5. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Sr no. Date Account Titles and Explanation Debit Credit
1. June 1, 2017
2. Oct. 1, 2018
3. Feb. 16, 2018
4. June 1, 2018
5. Nov. 2, 2018

Prepare the year-end adjusting entry required for items 1 through 7 on November 30. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. Round all amounts to the nearest dollar. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Sr no. Date Account Titles and Explanation Debit Credit
1. Nov. 30
2. Nov. 30
3. Nov. 30
4. Nov. 30
5. Nov. 30
6. Nov. 30
7. Nov. 30

Record the subsequent cash transactions in December for (1) the interest paid on December 1 (item 4), (2) the cheque received on December 4 (item 5), (3) the payroll paid on December 3 (item 6), and (4) the income tax paid on December 14 (item 7). (Round answers to 0 decimal places, e.g. 8,971. Credit account titles are automatically indented when the amount is entered. Do not indent manually.If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date Account Titles and Explanation Debit Credit
Dec. 1
Dec. 4
Dec. 3
Dec. 14
Answer #1
New Age Theatre Ltd.
Original Journal entries
No. Date Account Debit Credit Calculation
1 1-Jun-17 Vehicles $     80,700
Cash $     80,700
2 1-Oct-18 Cash $   170,200 =460*$370
Unearned Revenue $   170,200
3 16-Feb-18 Supplies $        2,370
Cash $        2,370
4 1-Jun-18 Cash $   105,300
Notes Payable $   105,300
5 2-Nov-18 Cash $           220
Unearned Revenue $           220
Adjusting Journal entries
1 30-Nov-18 Depreciation expense $     16,140 =80,700/5 years
Accumulated Depreciation $     16,140
2 30-Nov-18 Unearned Revenue $     63,825 =170,200/8*3
Ticket Revenue $     63,825 Oct to Nov - 3 plays
3 30-Nov-18 Supplies expense $        2,710 =1,010+2,370-670
Supplies $        2,710
4 30-Nov-18 Interest expense $           527 =105,300*6%*1/12
Interest Payable $           527
5 30-Nov-18 Unearned Revenue $           220
Rent Receivable $           300 =520-220
Rent Revenue $           520
6 30-Nov-18 Salaries expense $        6,257 =7,300/7*6
Salaries Payable $        6,257 Nov 25-30 payroll due
Sun-Fri
7 30-Nov-18 Income tax expense $        1,380
Income tax Payable $        1,380
Subsequent Journal entries
4 1-Dec-18 Interest Payable $           527
Cash $           527
5 4-Dec-18 Cash $           300
Rent Receivable $           300
6 3-Dec-18 Salaries Payable $        6,257
Salaries expense $        1,043 =7,300-6,257
Cash $        7,300
7 14-Dec-18 Income tax Payable $        1,380
Cash $        1,380